OpenStudy (anonymous):

Simplify square root of 3 multiplied by the fifth root of 3

OpenStudy (anonymous):

\[\sqrt{3} \sqrt[5]{3}\]

OpenStudy (anonymous):

\[\sqrt{3} \times \sqrt[5]{3} = 3^\frac{ 1 }{ 2 } * 3^\frac{ 1 }{ 5 } = 3^\frac{ 1 }{10}\] Or simple: \[\sqrt[10]{3}\] @jboy_52